OpenStudy (anonymous):

If f(x)=5x^3+3x^2-7x g(x)=3x^3-5x^2-2 h(x)=-9x^2+8 Find g(x)-f(x)

OpenStudy (pfenn1):

\[f(x) = 5x^3 +3x^2-7x\]\[g(x)=3x^3-5x^2 -2\]\[g(x)-f(x) = 5x^3 +3x^2 -7x - (3x^3 -5x^2-2)=\]\[(5-3)x^3 +(3+5)x^2-7x+2=2x^3+8x^2-7x+2\]

OpenStudy (anonymous):

Since it is g(x)-f(x), would it be -2x^3-2x^2-7x-2?

OpenStudy (pfenn1):

Did I do the math wrong above?

OpenStudy (anonymous):

I think we both did. We should have multiplied negative 1 to f(x) and we should have gotten -2x^3-8x^2+7x-2. Thanks for helping me work it through!

OpenStudy (pfenn1):

Isn't that what I did?\[g(x)-f(x) = 5x^3 +3x^2 -7x - (3x^3 -5x^2-2)=\]

OpenStudy (pfenn1):

\[g(x)-f(x) = (5x^3 +3x^2 -7x) - (3x^3 -5x^2-2)=\]

OpenStudy (pfenn1):

\[(5-3)x^3 +(3+5)x^2-7x+2=2x^3+8x^2-7x+2\]